OpenStudy (anonymous):

The ratio of a basketball player's completed free throws to attempted free throws is 4 to 7. If she completed 12 free throws, find how many free throws she attempted. a. 3 free throws c. 21 free throws b. 7 free throws d. 4 free throws

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

3.) It is recommended that there be at least 15.9 square feet of ground space in a garden for every newly planted shrub. A garden is 37.1 feet by 15 feet. Find the maximum number of shrubs the garden can accommodate. a. 185 shrubs c. 35 shrubs b. 12 shrubs d. 2 shrubs

OpenStudy (anonymous):

4.) Select the common unit rate you would use to solve a problem involving the speed of a vehicle. a. earnings per week c. kilometers per hour b. words per minute d. miles per gallon 5.)Jewelry is commonly weighed in carats. Five carats are equivalent to one gram. How many grams of gold are contained in a 24-carat gold chain? a. 120 g b. 4.8 g c. 5.0 g d. 100 g 6.)Jeremy's car went 300 miles (highway) on 10 gallons of gasoline. What was the average number of miles per gallon? a. 31 mpg c. 32 mpg b. 29 mpg d. 30 mpg 7.)In the Himalaya Mountains along the border of Tibet and Nepal, Mount Everest reaches a record height of 29,035 feet. How high is Mt. Everest in miles? Use the conversion equivalency of 1 mile = 5280 feet. a. 5.5 miles b. 5 miles c. 6 miles d. 3.5 miles 8.)To estimate the size of the grizzly bear population in a national park, rangers tagged and released 12 grizzly bears into the park. A few months later, 2 out of 21 grizzly bears sighted were tagged. Assuming the population of tagged bears in the population, estimate the number of grizzly bears in the population of this national park. a. 24 grizzly bears b. 126 grizzly bears c. 252 grizzly bears d. 22 grizzly bears

OpenStudy (nirmalnema):

the ans of 1st ques is 4 baskets out of 7 tries so for 12 baskets 7/4 *12 =21 tries

OpenStudy (anonymous):

can u help me answer the rest ?

OpenStudy (nirmalnema):

yup sure.. ans of 3 one.. is 35 shurbs total area of garden=37.1*15=556.5 area for 1shurb=15.9 so total no os plants = 556.5/15.9=35

OpenStudy (nirmalnema):

the ans os 4 is km per hr.. speed = distance / time km/ hr

OpenStudy (nirmalnema):

ans of 5 is 4.8 carat// 5 carat = 1 g 24 carat = 1/5 * 24=4.8

OpenStudy (nirmalnema):

ans of 6 is 30 gallons 300 miles for 10 gallons so 300/10 for one gallon = 30 miles

OpenStudy (nirmalnema):

7 ans.. 1 mile =5280 ft so for 29035 ft it is approx 5.5 miles

OpenStudy (nirmalnema):

ans 8 = 126 bears 2 out 21 os for 12 so total = 12*21 / 2 =126